OpenStudy (anonymous):

find the equilibrium point of S(x)=3x^2-3x+38 and D(x)= 56-6x

OpenStudy (amistre64):

when supply = demand

OpenStudy (amistre64):

it might be easier to subtract D(x) from S(x) and solve the quadratic

OpenStudy (amistre64):

S(x)=3x^2-3x + 38 - D(x)= (-6x + 56) --------------------- 3x^2 +3x -18 = 0 3(x^2 +x -6) = 0

OpenStudy (amistre64):

3(2) = 6 +x means +3-2 x=-(3,-2) = (-3,2)

OpenStudy (amistre64):

since the - is prolly a spurious result and undoable in real life; id go x=2

OpenStudy (amistre64):

since 44=44 id say its good